A monopolistic firm operates in two seperate markets. No trade is possible between market A and market B. The firm has calculated the demand functions for each market as follows;

Market A : P=15-Q Market B: P=11-Q

A company estimates its total cost function : TC=4Q

a)Quantity,total revenue and profit when company maximizes its profit and charges the same price in both markets

b)Quantity,total revenue and profit when the company charges different prices in each market and maximizes its total profit
1
Expert's answer
2017-05-25T11:37:10-0400
A) Pab=26-Q
TR=26Q-Q^2
MR=26-2Q
MC=4
26-2Q=4
Q=11
TR=26*11-11^2=165
Profit=165-44=121
b) TRa=15Q-Q^2
MRa=15-2Q
15-2q=4
Qa=5.5
TRa=52.25
TRb=11Q-Q^2
MRb=11-2Q
11-2Q=4
Qb=3.5
TRb=26.25
Qab=9
TRab=78.5
Profit=78.5-36=42.5